a belt runs on a wheel of radius 30cm. During the that wheel coasts uniformly to rest from an initial speed of 2.0revs/s, 25m of belt length passes over the wheel. find the deceleration of the wheel and the number of revolutions it turns while stopping
1
Expert's answer
2019-04-02T11:00:25-0400

Number of turns:


"n=\frac{l}{2\pi r}"

The angle covered during deceleration 


"\theta=2\pi n=\frac{l}{ r}"

The initial angular velocity


"\omega=2\pi (2)=4\pi"

Deceleration is


"\alpha=\frac{\omega^2}{2\theta}=\frac{16\pi^2}{2\frac{25}{0.3}}=0.95 rad/s^2"

The number of revolutions it turns while stopping


"n=\frac{25}{2\pi (0.3)}=13"